OpenStudy (sabrina1):

Suppose you have $5000 to invest. Calculate the annual interest rate r (in %) compounded continuously such that the invested money will triple in 20 years.

OpenStudy (sabrina1):

yes....

OpenStudy (sabrina1):

lolllll

OpenStudy (anonymous):

5000(r)^20

OpenStudy (anonymous):

sorry, 15000 = 5000(r)^20

OpenStudy (anonymous):

3 = r^20 log 3 = 20 log r r = 10^(log3 /20)

OpenStudy (sabrina1):

ok and how would i isolate for r?

OpenStudy (anonymous):

Do you know logs?

OpenStudy (mertsj):